OpenStudy (anonymous):

Solve the logarithmic equation. Be sure to reject any value that is not in the domain of the original logarithmic expressions. Give the exact answer. log6x2 = log6(5x + 36)

OpenStudy (anonymous):

A. (3/2) B. {9} C. Ø D. {9, -4}

OpenStudy (anonymous):

Your username/picture is awesome. *_*

OpenStudy (anonymous):

lol thanks ;-P

OpenStudy (anonymous):

You're welcome. (:

OpenStudy (anonymous):

log 6 + logx^2 = log6 + log(5x+36) logx^2 = log(5x+36) taking antilogs, x^2 = 5x + 36 What do you do after this

OpenStudy (anonymous):

\[x^2 -5x-36=0\\x^2+4x-9x-36\\ x(x+4)-9(x+4)\\(x-9)(x+4)\\ x=9 \ \ \ x=-4\]

OpenStudy (anonymous):

thank you ;-) @perurabo